OpenStudy (anonymous):

evaluate the function for relative extrema. evaluate the function for relative extrema. @Mathematics

OpenStudy (anonymous):

\[\sqrt{x^2+y^2}\]

OpenStudy (anonymous):

Is it just me or is the second partial test inconclusive

jimthompson5910 (jim_thompson5910):

so you want to find the relative extrema?

OpenStudy (anonymous):

yeah i did all the second partials and stuff but D=0 when i do it

OpenStudy (anonymous):

they say there is a relative min at (0,0,0)

jimthompson5910 (jim_thompson5910):

ok one sec

jimthompson5910 (jim_thompson5910):

that's correct

jimthompson5910 (jim_thompson5910):

that's actually a global min

OpenStudy (anonymous):

so does d=0 or something else

jimthompson5910 (jim_thompson5910):

but you have the right idea

jimthompson5910 (jim_thompson5910):

not sure what you're referring to with d

OpenStudy (anonymous):

not sure what it means... lol my teacher never said the name of it Fxx*Fyy-Fxy

OpenStudy (anonymous):

\[F_ {xx}*F_ {yy}-[F_ {xy}]^2=d\]

jimthompson5910 (jim_thompson5910):

ah found what you mean

OpenStudy (anonymous):

yeah my teacher was calling it distance for a while but then we got confused when he was using the distance formula lol

jimthompson5910 (jim_thompson5910):

D = fxx(a,b)*fyy(a,b) - {fxy(a,b)}², then 1) D > 0 and fxx > 0 => min 2) D > 0 and fxx < 0 => max 3) D < 0 => saddle point 4) D = 0 - indeterminate

jimthompson5910 (jim_thompson5910):

found here: http://www.physicsforums.com/showthread.php?t=195846

jimthompson5910 (jim_thompson5910):

i think he meant discriminant...but not sure tbh

OpenStudy (anonymous):

alright so when you did the second partials didn't you get 0*0-0^2=0

jimthompson5910 (jim_thompson5910):

let me differentiate real quick

OpenStudy (anonymous):

no matter what i always get x and y variables on top which will always = 0 but when you also use fxx,fxy,fyy(0,0,z) you get indeterminates

jimthompson5910 (jim_thompson5910):

yeah I'm getting division by zero errors

OpenStudy (anonymous):

well when you put it into form of a surface you get 0=x^2+y^2-z^2 which is going to be a point

jimthompson5910 (jim_thompson5910):

let me double check my work again, something might've gone wrong

OpenStudy (anonymous):

which doesn't work with a hyperboloid... so i'm confusedl ol

OpenStudy (anonymous):

if i trace it maybe...

jimthompson5910 (jim_thompson5910):

i think I know what's wrong

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

the theorem shown above only works with continuous partial second derivatives at (x,y)

OpenStudy (anonymous):

so how suppose i solve this?

OpenStudy (anonymous):

using the range?

jimthompson5910 (jim_thompson5910):

you're probably right in saying you have to work with 0=x^2+y^2-z^2

jimthompson5910 (jim_thompson5910):

just not sure how yet

OpenStudy (anonymous):

knowing that the range must be >=0 since it's square so the lowest it can be is 0?

jimthompson5910 (jim_thompson5910):

that's a good way to put it

OpenStudy (anonymous):

the thing is... that isn't a quadric surface ?

OpenStudy (anonymous):

the closest it is to is a hyperboloid of one sheet however that is not a hyperboloid of one sheet

OpenStudy (anonymous):

it looks sort of like a paraboloid

jimthompson5910 (jim_thompson5910):

paraboloid would be z = x^2+y^2 right?

jimthompson5910 (jim_thompson5910):

yeah it's a hyperboloid of one sheet

OpenStudy (anonymous):

not according to wolfram alpha?

jimthompson5910 (jim_thompson5910):

I'm looking at this http://en.wikipedia.org/wiki/Hyperboloid

jimthompson5910 (jim_thompson5910):

it's the first line

OpenStudy (anonymous):